OpenStudy (anonymous):

s(t)=)square root sign t+13... sorry I don't have a sign to do square roots.

OpenStudy (anonymous):

this is not really a question. you have \[s(t)=\sqrt{t+13}\]but what question is asked?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

s(-13) and s(13)

OpenStudy (anonymous):

evaluate the function?? I just don't know how to do it

OpenStudy (anonymous):

\[s(-13)=\sqrt{-13+13}\] \[\sqrt{0}=0\]

OpenStudy (anonymous):

\[s(13)=\sqrt{13+13}=\sqrt{26}\]

OpenStudy (anonymous):

thank u so much!